OpenStudy (anonymous):

Factor this expression as a difference of squares. 4-(2a+3b)^2

Parth (parthkohli):

We already have half of the answer. \( \color{Black}{\Rightarrow 2^2 - (2a + 3b)^2}\) \( \color{Black}{\Rightarrow (2 + 2a + 3b)(2 - (2a + 3b))}\) \( \color{Black}{\Rightarrow (2 + 2a + 3b)(2 - 2a - 3b) }\)

Parth (parthkohli):

Do you get it? @DeaneNicole

OpenStudy (anonymous):

To be honest.. No I'm completely lost @ParthKohli

Parth (parthkohli):

As you see - \((a + b)( a- b) = a^2 - b^2\) a is 2. b is 2a + 3b

Parth (parthkohli):

\( \color{Black}{\Rightarrow 2^2 - (2a + 3b) = (2 + 2a + 3b)(2 - (2a + 3b))}\) Exactly what I stated above

OpenStudy (anonymous):

Oh my gosh I'm so sorry, but, I'm still completely lost. D:

Parth (parthkohli):

I have a tutorial for this =) let me post it.

OpenStudy (anonymous):

Okay thank you so much!